OpenStudy (anonymous):

have quick question need to see if my answer is correct p(s)=2s^3-9s-10 P(1/2) the answer i got was 3/4 please someone work this and let me know

OpenStudy (anonymous):

your answer is wrong

OpenStudy (anonymous):

take it one step at a time. what is s^3 for s = 1/2?

OpenStudy (anonymous):

I need feedback to help you, tina

OpenStudy (anonymous):

you substitute s for 1/2

OpenStudy (anonymous):

sorry gave you the wrong equation P(s)=2s^3-2s-1

OpenStudy (anonymous):

your answer is still wrong. take it one step at a time. what is s^3 for s = 1/2?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

are you sure? s cubed is s times s times s.

OpenStudy (anonymous):

its not that complicated p(s)=2s^3-2s-1 p(1/2)=2(1/2)^3-2(1/2)-1

OpenStudy (anonymous):

right, so 2s^3 is 2 times 1/2 times 1/2 times 1/2 what is that equal to?

OpenStudy (anonymous):

2/8=1/4

OpenStudy (anonymous):

okay good. what is 2s?

OpenStudy (anonymous):

The value of 2s for s = 1/2 is?

OpenStudy (anonymous):

2/2=1 this is where i got messed up i think

OpenStudy (anonymous):

okay. so then your equation becomes (1/4)-1-1

OpenStudy (anonymous):

or (1/4)-2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so evaluate that. that value is your answer.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

how did you get that? post what you did here.

OpenStudy (anonymous):

1/4-2/1 2/4 1/2

OpenStudy (anonymous):

okay so are you telling me that if you take away 2 dollars from a quarter, you end up with 50 cents?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=52ZlXsFJULI go through that video. It will teach you how to how to add and subtract fractions.

OpenStudy (anonymous):

yeah cause i really hate fractions

OpenStudy (anonymous):

-1 3/4

OpenStudy (anonymous):

yes, that is right.

OpenStudy (anonymous):

go through some videos on how to add and subtract the fractions. The video I posted helps you think about fractions intuitively.

OpenStudy (anonymous):

ok so then that is the final answer, i will i get them mixed up all the time

OpenStudy (anonymous):

okay goodluck.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np